OpenStudy (anonymous):

Is the solution to the radical equation the square root of 6 x equals x x = 0, x = 6 or both?

OpenStudy (anonymous):

sqrt (6x) = x, (6x) = x^2 x^2 - 6x = 0 (x)(x-6) = 0. x=0 and x-6 = 0 Both